OpenStudy (anonymous):

3 fifth root x plus 2 cubed + 3 = 27

OpenStudy (anonymous):

((3/5)^x) + 2^3 = 27

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

Yes

OpenStudy (bibby):

\(3 \sqrt[5]{x} + 2^3+ 3 = 27\)?

OpenStudy (anonymous):

oh that might be it...

OpenStudy (anonymous):

after a few revisions I think (3/5)^x is the right way to do it. cause root isn't square root root is like ^ like 5^2 or something